OpenStudy (abbycross167):

Can someone please help me answer this question? Graph the following systems. Draw one line in yellow and one line in blue. Indicate the solution to the system ON the GRAPH. x - y = 0 3x + y = -4

RhondaSommer (rhondasommer):

https://www.desmos.com/calculator

RhondaSommer (rhondasommer):

does that help?

OpenStudy (abbycross167):

Somewhat, my teacher said that I need to show my solution. I did the graph right

RhondaSommer (rhondasommer):

y=mx+b so \[x-y=0\] subtract x from both sides \[-y=x\] divide both sides by negative 1\[y=-x\]

RhondaSommer (rhondasommer):

y=mx+b so \[3x + y = -4\] subtract 3x from both sides so \[y=-3x-4\]

RhondaSommer (rhondasommer):

you can graph both of those :)

RhondaSommer (rhondasommer):

make more sense?

RhondaSommer (rhondasommer):

\[y=x\] sorry mistype on me :/

OpenStudy (abbycross167):

yes ma'am/sir but is that the solution that I need to graph? And where do I put y=x? @RhondaSommer

RhondaSommer (rhondasommer):

\[y=mx+b\] so \[x−y=0\] subtract x from both sides \[−y=-x\] divide both sides by negative 1 \[y=x\]

OpenStudy (abbycross167):

oh ok

RhondaSommer (rhondasommer):

im genderfluid; it doesnt matter what you call me

RhondaSommer (rhondasommer):

:)

OpenStudy (abbycross167):

oh ok, but that is the solution that my teacher was talking about?

RhondaSommer (rhondasommer):

i have no idea; we can get what x and y equals. we could put it in another equation...youd have to ask

RhondaSommer (rhondasommer):

did you want to solve for x and y?